Danyourman1:

Can someone help me with this, I don't get it. Set up a system of equations and solve using substitution or elimination. You must show all work. You have a stack of bills with $1 bills and $5 bills in it. You have 43 total bills equaling $99. How many of each type of bill do you have? Number of $1 bills: _________________ Number of $5 bills: __________________

Vocaloid:

Let’s let o = # of one dollar bills and f = # of five dollar bills We have 43 bills in total so: o + f = 43 The bills are also worth $99 in total. If you have (o) one dollar bills, they are worth 1*o or simply o dollars. Similarly, f five dollar bills are worth 5f. So o + 5f = 99 So your system is now o + f = 43 o + 5f = 99 Can you try solving the system now? Elimination looks like a good choice due to how it’s set up

Danyourman1:

thank you